OpenStudy (anonymous):

what is the sum of the sequence?1+3+5+7...+1001

OpenStudy (anonymous):

Let n be no. of terms. It is an A.P. a=1,c.d d=3-1=2 l=1001 l=a+(n-1)d 1001=1+(n-1)2 1000/2=n-1 n=500+1=501 \[sn=\frac{ n \left( a+l \right) }{ 2 }=\frac{ 501\left( 1+1001 \right) }{2 }=501*501=?\] calculate.